OpenStudy (anonymous):

Simplify 3(x + 2) + 4(x - 5)

OpenStudy (linknissan):

3x+3*2+4x-4*5=0 3x+6+4x-20=0 7x-12=0 x=12/7

OpenStudy (anonymous):

@linknissan -20+6 is not equal to 12...

OpenStudy (linknissan):

oh srry 14

OpenStudy (anonymous):

@ryantorres got it..??

OpenStudy (anonymous):

yes , thank you so much !